OpenStudy (anonymous):

solve 5^(x + 5) = 9^x ?

OpenStudy (anonymous):

\[\large 5^{x+5}=9^x ?\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

do u know the properties of log? and can u use calculator?

OpenStudy (anonymous):

no i know how to solve it. I just don't have a calculator at the moment.

OpenStudy (anonymous):

so u just want the answer?

OpenStudy (anonymous):

yes plz

OpenStudy (anonymous):

take log \[\huge (x+5) \log 5 =x \log 9 \\ \huge x(\log 5- \log 9)=-5 \log 5 \\ \huge x=\frac{5 \log 5}{\log 9-\log 5}=13.69\]

OpenStudy (anonymous):

ok thankyou so much!!